# FACET: Teacher-Centred LLM-Based Multi-Agent Systems— Towards Personalized Educational Worksheets

### v1

@@ Appendix C Appendix C
1 | Clear didactical progression, from simple to complex, from concrete to abstract, and goals are transparent. |
2 | Mostly clear structure and goal transparency, useful guidance for different levels. |
3 | No visible adaptation to student perspective. |
4 | Structure is hard to follow or overly complex without didactical necessity. |
5 | Confusing or misleading structure; goals unclear; no consolidation or even contradictory instructions. |
6 | No worksheet or completely unusable structure. |
@@
1 | Task is clearly formulated, logically sequenced, easy to understand for the target group. |
2 | Mostly clear with only minor ambiguities; task goal is still understandable. |
3 | No special consideration for clarity; some unclear formulations. |
4 | Overcomplicated or verbose formulation that makes the task harder than necessary. |
5 | Formulation is misleading or includes wrong cues or terminology. |
6 | Task is missing or incomprehensible. |
@@
1 | Highly original or open-ended task allowing for multiple perspectives or solutions; stimulates creative thinking. |
2 | Some creative elements or opportunities for individual approaches. |
3 | Standard task, no special creative value. |
4 | Task forces pseudo-creative framing without real value or makes task harder through unnecessary open-endedness. |
5 | ‘Creative’ elements are misleading or inappropriate for the learning goal. |
6 | No creative aspects or task is missing. |
@@
1 | Perfectly adapted to the learners’ prior knowledge, interests, and level of difficulty; starts from the concrete/known. |
2 | Well adapted, minor mismatches with regard to content or level. |
3 | No effort made to adapt to learner level (e.g., too generic or average). |
4 | Adaptation attempt made, but misjudges learner needs or overcomplicates. |
5 | Task is developmentally inappropriate or contains false assumptions about learner ability. |
6 | Task not completed or fully unsuitable for intended learners. |

### v2

@@ Appendix C Appendix C
| 1 | Clear didactical progression, from simple to complex, from concrete to abstract, and goals are transparent. |
| 2 | Mostly clear structure and goal transparency, useful guidance for different levels. |
| 3 | No visible adaptation to student perspective. |
| 4 | Structure is hard to follow or overly complex without didactical necessity. |
| 5 | Confusing or misleading structure; goals unclear; no consolidation or even contradictory instructions. |
| 6 | No worksheet or completely unusable structure. |
@@
| 1 | Task is clearly formulated, logically sequenced, easy to understand for the target group. |
| 2 | Mostly clear with only minor ambiguities; task goal is still understandable. |
| 3 | No special consideration for clarity; some unclear formulations. |
| 4 | Overcomplicated or verbose formulation that makes the task harder than necessary. |
| 5 | Formulation is misleading or includes wrong cues or terminology. |
| 6 | Task is missing or incomprehensible. |
@@
| 1 | Highly original or open-ended task allowing for multiple perspectives or solutions; stimulates creative thinking. |
| 2 | Some creative elements or opportunities for individual approaches. |
| 3 | Standard task, no special creative value. |
| 4 | Task forces pseudo-creative framing without real value or makes task harder through unnecessary open-endedness. |
| 5 | ‘Creative’ elements are misleading or inappropriate for the learning goal. |
| 6 | No creative aspects or task is missing. |
@@
| 1 | Perfectly adapted to the learners’ prior knowledge, interests, and level of difficulty; starts from the concrete/known. |
| 2 | Well adapted, minor mismatches with regard to content or level. |
| 3 | No effort made to adapt to learner level (e.g., too generic or average). |
| 4 | Adaptation attempt made, but misjudges learner needs or overcomplicates. |
| 5 | Task is developmentally inappropriate or contains false assumptions about learner ability. |
| 6 | Task not completed or fully unsuitable for intended learners. |